# DAG-aware Transformer for Causal Effect Estimation

### v1

@@ Abstract
Causal inference is a critical task across fields such as healthcare, economics, and the social sciences. While recent advances in machine learning, especially those based on the deep-learning architectures, have shown potential in estimating causal effects, existing approaches often fall short in handling complex causal structures and lack adaptability across various causal scenarios. In this paper, we present a novel transformer-based method for causal inference that overcomes these challenges. The core innovation of our model lies in its integration of causal Directed Acyclic Graphs (DAGs) directly into the attention mechanism, enabling it to accurately model the underlying causal structure. This allows for flexible estimation of both average treatment effects (ATE) and conditional average treatment effects (CATE). Extensive experiments on both synthetic and real-world datasets demonstrate that our approach surpasses existing methods in estimating causal effects across a wide range of scenarios. The flexibility and robustness of our model make it a valuable tool for researchers and practitioners tackling complex causal inference problems.
@@ 1 Introduction
The estimation of Average Treatment Effect (ATE) and Conditional Average Treatment Effect (CATE) plays a pivotal role across various disciplines, significantly impacting decision-making processes and policy formulation. In medicine, these estimations guide treatment selections and personalized healthcare strategies (Hernán and Robins, 2024; Glass et al., 2013; Wager and Athey, 2018). Within the realm of public policy, they inform the design and evaluation of interventions, from education reforms to social welfare programs (Imbens and Rubin, 2015; Hill, 2011). In economics, ATE and CATE estimations are crucial for understanding the impacts of economic policies, labor market interventions, and consumer behavior (Angrist and Pischke, 2008; Heckman and Vytlacil, 2007).
@@
A fundamental challenge in this field lies in the correct specification of propensity score and outcome models, particularly when employing methods such as Inverse Probability of Treatment Weighting (IPTW) and Doubly-Robust Estimator (or Augmented IPW) to control for confounding factors (Hernán and Robins, 2024; Robins et al., 1994; Bang and Robins, 2005). These methods, while powerful, are sensitive to model misspecification, which can lead to biased estimates and potentially misleading conclusions (Kang and Schafer, 2007; Funk et al., 2011). The complexity of real-world scenarios, characterized by high-dimensional data and complex causal relationships, further exacerbates this challenge, necessitating more sophisticated and robust approaches to causal inference (Chernozhukov et al., 2018; Wager and Athey, 2018).
The integration of machine learning (ML) methods into causal inference has opened new avenues for addressing complex causal relationships in high-dimensional settings. Athey and Imbens (2016) introduced causal trees and forests, adapting random forest algorithms to estimate heterogeneous treatment effects with valid statistical inference. Expanding on this, Wager and Athey (2018) developed generalized random forests, extending forest-based methods to a broader class of causal parameters. These approaches have shown promise in settings with many covariates and potential treatment effect heterogeneity. Concurrently, Chernozhukov et al. (2018) proposed the double machine learning framework, combining flexible ML methods with orthogonalization techniques to achieve valid inference on treatment effects in high-dimensional settings.
@@
Deep learning methods have also made significant inroads in causal inference, offering powerful tools for modeling complex relationships. Shalit et al. (2017) introduced representation learning techniques for estimating individual treatment effects, using neural networks to learn balanced representations of covariates, addressing the fundamental problem of unobserved counterfactuals. The emergence of graph neural networks (GNNs) has further expanded the possibilities in causal inference, particularly for networked data. Ma and Tresp (2020) demonstrated how GNN-based approaches can estimate heterogeneous treatment effects in the presence of spillover effects, capturing complex dependencies in networked experiments. These deep learning methods offer new ways to handle the challenges of high-dimensional data and complex causal structures in modern causal inference problems. Recent work has explored transformer architectures for causal inference tasks. Melnychuk et al. (2022) introduced the Causal Transformer for estimating counterfactual outcomes over time, effectively capturing long-range dependencies in longitudinal data. Zhang et al. (2023b) proposed TransTEE, a transformer-based model for Heterogeneous Treatment Effect (HTE) estimation that handles various types of treatments. Zhang et al. (2023a) developed Causal Inference with Attention (CInA), enabling zero-shot causal inference on unseen tasks with new data.
@@
To address these limitations, we propose a novel approach that harnesses the power of transformer models while explicitly incorporating causal structure through a DAG-aware attention mechanism. Our method enables the estimation of crucial causal quantities including the propensity score model $P(A|\mathbf{X})$, the outcome regression model $P(Y|A,\mathbf{X})$, and the bridge function $h(A,W,X)$. Here, $A$ represents the treatment, $\mathbf{X}$ denotes observed confounders, $Y$ is the outcome, and $W$ serves as a proxy for the outcome in scenarios with unmeasured confounding.
This approach allows for seamless integration of these estimated models into IPTW, doubly robust estimators, and proximal inference methods. By doing so, our work bridges the gap between cutting-edge machine learning techniques and classical causal inference methods, offering a more comprehensive framework for causal analysis in complex, real-world scenarios.
@@
  • Development of a DAG-aware transformer model that explicitly incorporates causal structure into the attention mechanism, allowing for more accurate modeling of causal relationships.
  • Concurrent estimation of propensity score and outcome models within a single, unified framework, improving efficiency and potentially reducing bias in causal effect estimation.
  • Seamless integration of the proposed method with established causal inference techniques such as IPTW, doubly robust estimation and proximal inference, enhancing their performance in complex, high-dimensional settings.
  • Empirical evaluation of the proposed method on both simulated and real-world datasets, demonstrating its effectiveness in estimating causal effects across various scenarios.
@@ 2.1 ATE and CATE
2.2 Confounding Control Methods
In causal inference, several methods have been developed to control for observed confounding and estimate treatment effects. While we explored multiple approaches, our paper focuses primarily on two methods: Inverse Probability of Treatment Weighting (IPTW) and Augmented Inverse Probability Weighting (AIPW), a form of Doubly Robust estimator (Hernán and Robins, 2024). We also considered proximal inference (Tchetgen et al., 2024) to incorporate unmeasured confounding; this method is described in Appendix C.
@@ 2.2 Confounding Control Methods
    $\tau_{AIPW}=\mathbb{E}\Big{[}\big{(}\mu(1,X)+\frac{A}{\pi(X)}(Y-\mu(1,X))\big{)}-\big{(}\mu(0,X)+\frac{1-A}{1-\pi(X)}(Y-\mu(0,X))\big{)}\Big{]}$ (2)
@@ 3 Methodology
We propose a novel DAG-aware Transformer model for causal effect estimation that explicitly incorporates causal structure into the attention mechanism. Here we describe settings assuming no unmeasured confounding, and we extend our model to accommodate unmeasured confounding using proximal inference in Appendix C. Given a dataset of $N$ observations, we define input nodes $\mathcal{X}$, which include the treatment $A$, observed confounding variables $\mathbf{X}$, and outcome $Y$. The output nodes are $\hat{A}$ and $\hat{Y}$. Our objective is to estimate both the Average Treatment Effect (ATE) and the Conditional Average Treatment Effect (CATE).
@@ 3.1 DAG-aware Transformer Architecture
Adjacency MatrixDAG-aware AttentionOutput Nodes ($\hat{A}$, $\hat{Y}$)
Figure 1: Architecture of the DAG-aware Transformer model. The Transformer Encoder incorporates the DAG-aware attention mechanism (highlighted with dashed lines), which utilizes the causal structure represented by the DAG. The adjacency matrix derived from the causal DAG informs the DAG-aware attention computation. For simplicity, layer normalization and feed-forward networks within the Transformer Encoder are not shown.
This mask ensures attention flows only along causal pathways and allows self-attention for each node. Our key innovation lies in incorporating the causal structure into the attention mechanism. In each multi-head attention layer, we compute attention scores $\mathbf{A}=\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{E}}$ and apply the DAG-based mask $\mathbf{A}^{mask}=\mathbf{A}+\mathbf{M}\cdot(-\infty)$, where $\mathbf{Q},\mathbf{K}\in\mathbb{R}^{N\times D\times E}$ are the query and key matrices, and $E$ is the embedding dimension. This operation effectively sets attention scores to zero (after softmax) for node pairs not causally linked in the DAG.
The masked attention scores are then normalized using softmax and used to compute the output:
$\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}(\mathbf{A}^{mask})\mathbf{V}$ (4)
where $\mathbf{V}$ is the value matrix.
A key feature of our approach is the simultaneous estimation of both the propensity score model and the outcome regression model within the same architecture. The propensity score model $\hat{\pi}(x)=P(A=1|X=x)$ is estimated through the output head for node $A$. Given input features $X$, the model outputs a probability distribution over treatment assignment. Simultaneously, the outcome model $\hat{\mu}(a,x)=E[Y|A=a,X=x]$ is estimated through the output head for node $Y$, considering both $X$ and $A$. These estimates are then plugged into the IPTW and AIPW estimators to compute the ATE and CATE.
@@ 3.2 Model Training and Objective Function
Building upon the work of Saito and Yasui (2020), we train our DAG-aware Transformer model in an end-to-end manner. The model learns representations $\Phi(\mathbf{X})$ for the treatment and $h(\Phi(\mathbf{X}),A)$ for the outcome simultaneously. We employ the Adam optimizer with weight decay regularization (Loshchilov and Hutter, 2019) for training.
Our objective function combines a weighted empirical risk term with a regularization term based on the Integral Probability Metric (IPM). This formulation allows us to balance between fitting the observed data and ensuring that the learned representations ($\Phi(\mathbf{X})$) are similar for the treated and control groups. The objective function is defined as:
$h^{\ast}=\min_{h}\Bigg{[}\underbrace{\frac{1}{n}\sum_{i=1}^{n}w^{\prime}_{a}(x_{i})L(h(\Phi(x_{i}),a_{i}),y_{i})}_{\text{empirical weighted risk}}+\alpha\underbrace{\text{IPM}_{G}\left(\{\Phi(x_{i})\}_{i:a_{i}=0},\{\Phi(x_{i})\}_{i:a_{i}=1}\right)}_{\text{distributional distance}}\Bigg{]}$ (5)
In this equation, $w^{\prime}_{a}(x_{i})=\frac{w_{a}(x_{i})}{2}(\frac{a_{i}}{\hat{\pi}_{1}}+\frac{1-a_{i}}{\hat{\pi}_{0}})$ represents the sample weight, where $\hat{\pi}_{a}=\frac{1}{n}\sum_{i=1}^{n}I\{a_{i}=a\}$ is the empirical probability of treatment assignment. The stabilizing weight $w_{a}(x_{i})$ is defined as $w_{a}(x_{i})=\frac{a(1-2e(x_{i}))+e(x_{i})^{2}}{e(x_{i})(1-e(x_{i}))}$, with $e(x_{i})$ being the propensity score. The trade-off hyperparameter $\alpha$ controls the balance between the empirical risk and the distributional distance. For the Integral Probability Metric $\text{IPM}_{G}$, we use the Wasserstein distance in our experiments. The empirical weighted risk term ensures that our model fits the observed data well, while the distributional distance term encourages similarity between the representations of the treated and control groups. It’s important to note that while the IPM term encourages balancing between the treated and control groups, excessive balancing could potentially increase the Representation-Induced Confounding Bias (RICB) (Melnychuk et al., 2024). This occurs because strong balancing might lead to a loss of information about confounders in the learned representations. Therefore, careful tuning of the hyperparameter $\alpha$ is crucial to achieve an optimal trade-off between balancing and preserving important confounding information.

### v2

@@ Abstract
Causal inference is a critical task across fields such as healthcare, economics, and the social sciences. While recent advances in machine learning, especially those based on the deep-learning architectures, have shown potential in estimating causal effects, existing approaches often fall short in handling complex causal structures and lack adaptability across various causal scenarios. In this paper, we present a novel transformer-based method for causal inference that overcomes these challenges. The core innovation of our model lies in its integration of causal Directed Acyclic Graphs (DAGs) directly into the attention mechanism, enabling it to accurately model the underlying causal structure. This allows for flexible estimation of both average treatment effects (ATE) and conditional average treatment effects (CATE). Extensive experiments on both synthetic and real-world datasets demonstrate that our approach surpasses existing methods in estimating causal effects across a wide range of scenarios. The flexibility and robustness of our model make it a valuable tool for researchers and practitioners tackling complex causal inference problems. Our code is available at https://github.com/ManqingLiu/DAGawareTransformer.
Machine Learning, causality, DAG, transformer, ICML
@@ 1 Introduction
The estimation of Average Treatment Effect (ATE) and Conditional Average Treatment Effect (CATE) plays a pivotal role across various disciplines, significantly impacting decision-making processes and policy formulation. In medicine, these estimations guide treatment selections and personalized healthcare strategies (Hernán & Robins, 2024; Glass et al., 2013; Wager & Athey, 2018). Within the realm of public policy, they inform the design and evaluation of interventions, from education reforms to social welfare programs (Imbens & Rubin, 2015; Hill, 2011). In economics, ATE and CATE estimations are crucial for understanding the impacts of economic policies, labor market interventions, and consumer behavior (Angrist & Pischke, 2008; Heckman & Vytlacil, 2007).
@@
A fundamental challenge in this field lies in the correct specification of propensity score and outcome models, particularly when employing methods such as Inverse Probability of Treatment Weighting (IPTW) and Doubly-Robust Estimator (or Augmented IPW) to control for confounding factors (Hernán & Robins, 2024; Robins et al., 1994; Bang & Robins, 2005). These methods, while powerful, are sensitive to model misspecification, which can lead to biased estimates and potentially misleading conclusions (Kang & Schafer, 2007; Funk et al., 2011). The complexity of real-world scenarios, characterized by high-dimensional data and complex causal relationships, further exacerbates this challenge, necessitating more sophisticated and robust approaches to causal inference (Chernozhukov et al., 2018; Wager & Athey, 2018).
@@
The integration of machine learning (ML) methods into causal inference has opened new avenues for addressing complex causal relationships in high-dimensional settings. Athey and Imbens (Athey & Imbens, 2016) introduced causal trees and forests, adapting random forest algorithms to estimate heterogeneous treatment effects with valid statistical inference. Expanding on this, Wager and Athey (Wager & Athey, 2018) developed generalized random forests, extending forest-based methods to a broader class of causal parameters. These approaches have shown promise in settings with many covariates and potential treatment effect heterogeneity. Concurrently, Chernozhukov et al. (Chernozhukov et al., 2018) proposed the double machine learning framework, combining flexible ML methods with orthogonalization techniques to achieve valid inference on treatment effects in high-dimensional settings. Deep learning methods have also made significant inroads in causal inference, offering powerful tools for modeling complex relationships. Shalit et al. (Shalit et al., 2017) introduced representation learning techniques for estimating individual treatment effects, using neural networks to learn balanced representations of covariates, addressing the fundamental problem of unobserved counterfactuals. The emergence of graph neural networks (GNNs) has further expanded the possibilities in causal inference, particularly for networked data. Ma et al. (Ma & Tresp, 2020) demonstrated how GNN-based approaches can estimate heterogeneous treatment effects in the presence of spillover effects, capturing complex dependencies in networked experiments. Recent work has also explored transformer architectures for causal inference tasks. Melnychuk et al. (Melnychuk et al., 2022) introduced the Causal Transformer for estimating counterfactual outcomes over time, effectively capturing long-range dependencies in longitudinal data. Zhang et al. (Zhang et al., 2023b) proposed TransTEE, a transformer-based model for Heterogeneous Treatment Effect (HTE) estimation that handles various types of treatments. Zhang et al. (Zhang et al., 2023a) developed Causal Inference with Attention (CInA), enabling zero-shot causal inference on unseen tasks with new data. These deep learning methods offer new ways to handle the challenges of high-dimensional data and complex causal structures in modern causal inference problems.
@@
To address these limitations, we propose a novel approach that harnesses the power of transformer models while explicitly incorporating causal structure through a DAG-aware attention mechanism. Our method enables the estimation of crucial causal quantities including the propensity score model $P(A|\mathbf{X})$, the outcome regression model $P(Y|A,\mathbf{X})$, and the bridge function $h(A,W,X)$. Here, $A$ represents the treatment, $\mathbf{X}$ denotes observed confounders, $Y$ is the outcome, and $W$ serves as a proxy for the outcome in scenarios with unmeasured confounding. This approach allows for seamless integration of these estimated models into IPTW, doubly robust estimators, and proximal inference methods. By doing so, our work bridges the gap between cutting-edge machine learning techniques and classical causal inference methods, offering a more comprehensive framework for causal analysis in complex, real-world scenarios.
@@
  • Development of a DAG-aware Transformer model that explicitly incorporates causal structure into the attention mechanism, allowing for more accurate modeling of causal relationships in various estimation frameworks including G-formula, IPW, and AIPW.
  • Demonstration of the model’s versatility in both joint and separate training paradigms for propensity score and outcome models, revealing context-dependent benefits: joint training excels in ATE estimation, while separate training shows advantages in CATE estimation.
  • Empirical evaluation of the proposed method on diverse datasets (Lalonde-CPS, Lalonde-PSID, ACIC, and Demand), showcasing its superior performance in estimating both average and conditional average treatment effects across various scenarios and sample sizes.
  • Extension of the DAG-aware Transformer to proximal causal inference, demonstrating significant improvements over existing methods, particularly in the NMMR-U framework across different training set sizes.
@@ 2.1 ATE and CATE
2.2 Confounding Control Methods assuming Unconfoundedness
In causal inference, several methods have been developed to control for observed confounding and estimate treatment effects. Our paper focuses primarily on three methods: Standardization (G-formula), Inverse Probability of Treatment Weighting (IPTW) and Augmented Inverse Probability Weighting (AIPW), a form of Doubly Robust estimator (Hernán & Robins, 2024).
@@ 2.2 Confounding Control Methods assuming Unconfoundedness
    Standardization (G-formula): Standardization, also known as the G-formula, estimates the ATE by modeling the outcome as a function of treatment and confounders. It then averages over the confounder distribution to estimate the population-level effect. The ATE is estimated as:
    $\tau_{G}=\mathbb{E}_{X}[\mathbb{E}[Y|A=1,X]-\mathbb{E}[Y|A=0,X]]$ (1)
    where $\mu(a,X)=\mathbb{E}[Y|A=a,X]$ is the conditional expectation of the outcome given treatment $a$ and confounders $X$. This method is effective when the outcome model is correctly specified.
  2. 2.
@@
    $\tau_{AIPW}=\mathbb{E}\Big{[}$ $\big{(}\mu(1,X)+\frac{A}{\pi(X)}(Y-\mu(1,X))\big{)}$ (3)
    $-\big{(}\mu(0,X)+\frac{1-A}{1-\pi(X)}(Y-\mu(0,X))\big{)}\Big{]}$
@@
2.3 Proximal Inference
In proximal inference (Tchetgen et al., 2024), we aim to estimate the expected potential outcome $\mathbb{E}[Y^{a}]$ for each treatment level $a$, in the presence of unobserved confounders $U$, given a set of proxies $(W,Z)$ and observed confounders $X$. The key assumptions are:
Assumption 1.
Assumption 2.
Assumption 3.
[…]
The expected potential outcomes are given by:
$\mathbb{E}[Y^{a}]=\mathbb{E}_{W,X}[h(a,W,X)]$ (5)
@@ 3 Methodology
We propose a novel DAG-aware Transformer model for causal effect estimation that explicitly incorporates causal structure into the attention mechanism. Our approach is flexible and can accommodate various causal scenarios, including those with or without unmeasured confounding.
Given a dataset of $N$ observations, we define a set of possible input nodes. These include: $A$, the treatment variable; $\mathbf{X}$, the observed confounding variables; $U$, representing unmeasured confounding variables; $Y$, the outcome variable; $Z$, the proxy variable for treatment; and $W$, the proxy variable for outcome. The specific combination of input nodes used depends on the causal structure being modeled. The output nodes of our model vary based on the estimation method employed:
  • For standardization or proximal inference: $\hat{Y}$ (estimated outcome), which is $\hat{\mu}(a,X)$ for standardization or $\hat{h}(a,W,X)$ for proximal inference.
  • For Inverse Probability of Treatment Weighting (IPTW): $\hat{A}=\hat{\pi}(X)$ (estimated propensity score)
  • For Augmented Inverse Probability Weighting (AIPW): Both $\hat{A}$ and $\hat{Y}$
This flexible framework allows our DAG-aware Transformer to adapt to different causal inference scenarios and estimation techniques while maintaining its core structure.
After estimating $\hat{A}$ and $\hat{Y}$, we can plug these values into the corresponding equations to estimate the Average Treatment Effect (ATE) or Conditional Average Treatment Effect (CATE). For instance:
  • For standardization: ATE = $\mathbb{E}_{X}[\hat{\mu}(1,X)-\hat{\mu}(0,X)]$
  • For IPTW: ATE = $\mathbb{E}[\frac{AY}{\hat{\pi}(X)}-\frac{(1-A)Y}{1-\hat{\pi}(X)}]$
  • For AIPW: ATE = $\mathbb{E}[(\hat{\mu}(1,X)-\hat{\mu}(0,X))+\frac{A(Y-\hat{\mu}(1,X))}{\hat{\pi}(X)}-\frac{(1-A)(Y-\hat{\mu}(0,X))}{1-\hat{\pi}(X)}]$
  • For proximal inference: ATE = $\mathbb{E}[\hat{h}(1,W,X)-\hat{h}(0,W,X)]$
For CATE estimation, we can condition on specific values of $X$ in these equations. This approach allows us to estimate both population-level and subgroup-level causal effects using our DAG-aware Transformer model.
@@ 3.1 DAG-aware Transformer Architecture
This mask ensures attention flows only along causal pathways and allows self-attention for each node.
Our key innovation lies in incorporating the causal structure into the attention mechanism. In each multi-head attention layer, we compute attention scores $\mathbf{A}=\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{E}}$ and apply the DAG-based mask $\mathbf{A}^{mask}=\mathbf{A}+\mathbf{M}\cdot(-\infty)$, where $\mathbf{Q},\mathbf{K}\in\mathbb{R}^{N\times D\times E}$ are the query and key matrices, and $E$ is the embedding dimension. This operation effectively sets attention scores to zero (after softmax) for node pairs not causally linked in the DAG.
The masked attention scores are then normalized using softmax and used to compute the output: $\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}(\mathbf{A}^{mask})\mathbf{V}$ where $\mathbf{V}$ is the value matrix.
After the transformer encoder processes the input, we introduce a novel component to our architecture. We concatenate the output from the encoder (weighted by $alpha$) with the raw input before passing it to a Multi-Layer Perceptron (MLP). This approach is inspired by residual connections used in ResNet architectures (He et al., 2016). The weighted output is computed as follows: $\mathbf{Z}=\alpha\cdot\mathbf{H}+\mathbf{X}$ where $\mathbf{H}$ is the output from the Transformer Encoder, $\mathbf{X}$ is the raw input, and $\alpha$ is a hyper parameter.
The rationale behind this concatenation is to mitigate potential loss of confounding information that might occur when using the encoder alone. We found that relying solely on the encoder output could sometimes result in identical predictions of $\hat{Y}$ regardless of the treatment value, indicating a loss of important confounding information. By incorporating the raw input directly, we ensure that all relevant information is preserved and utilized in the final prediction stage.
The weighted output $\mathbf{Z}$ is then fed into an MLP, which produces the final output: $\hat{A}/\hat{Y}=\text{MLP}(\mathbf{Z})$
This architecture allows our model to leverage both the complex representations learned by the transformer encoder and the raw input features, leading to more robust and accurate causal effect estimations.
@@
Adjacency MatrixDAG-aware AttentionMLPOutput Nodes ($\hat{A}$, $\hat{Y}$) Concatenation of encoder output and raw input
Figure 1: Architecture of the DAG-aware Transformer model. The Transformer Encoder incorporates the DAG-aware attention mechanism (highlighted with dashed lines), which utilizes the causal structure represented by the DAG. The adjacency matrix derived from the causal DAG informs the DAG-aware attention computation. The model combines the output from the transformer encoder with the raw input through a weighted average, which is then processed by an MLP to produce the final output. For simplicity, layer normalization and feed-forward networks within the transformer encoder are not shown.
@@ 3.2 Model Training and Objective Function
Our model employs different loss functions depending on the causal inference method used. We present the loss functions for standardization (G-formula), Inverse Probability of Treatment Weighting (IPTW), Augmented Inverse Probability Weighting (AIPW), and proximal inference. Here we assume $Y$ (outcome) is continuous, and $A$ (treatment) is a binary variable.
3.2.1 Standardization
For standardization, we use Mean Squared Error (MSE) loss for maximum likelihood estimation:
$\mathcal{L}_{G}=\text{MSE}(\hat{Y},Y)=\frac{1}{n}\sum_{i=1}^{n}(\hat{Y}_{i}-Y_{i})^{2}$ (7)
where $\hat{Y}$ are the model outputs and $Y$ are the true labels.
3.2.2 Inverse Probability of Treatment Weighting (IPTW)
For IPTW, we use Binary Cross Entropy (BCE) loss for treatment/propensity score estimation:
$\mathcal{L}_{\text{IPTW}}=\text{BCE}(\hat{A},A)=$ (8)
$-\frac{1}{n}\sum_{i=1}^{n}[A_{i}\log(\hat{A}_{i})+(1-A_{i})\log(1-\hat{A}_{i})]$
where $\hat{A}$ are the model outputs (estimated propensity scores) and $A$ are the true treatment assignments.
3.2.3 Augmented Inverse Probability Weighting (AIPW)
For AIPW, we combine MSE loss for outcome prediction and BCE loss for treatment assignment:
$\mathcal{L}_{\text{AIPW}}=\frac{1}{2}(\text{MSE}(\hat{Y},Y)+\text{BCE}(\hat{A},A))$ (9)
3.2.4 Proximal Inference
For proximal inference, following the work of (Kompa et al., 2022), we introduce two variants: NMMR-U and NMMR-V, based on U-statistics and V-statistics respectively. The empirical risk $\hat{R}_{k,n}$ given data $\mathcal{D}=\{(a_{i},w_{i},x_{i},y_{i},z_{i})\}_{i=1}^{N}$ can be written as:
$\hat{R}_{k,U,n}(h)=\frac{1}{n(n-1)}\sum_{i,j=1,i\neq j}^{n}(y_{i}-h_{i})(y_{j}-h_{j})k_{ij}$ (10)
$\hat{R}_{k,V,n}(h)=\frac{1}{n^{2}}\sum_{i,j=1}^{n}(y_{i}-h_{i})(y_{j}-h_{j})k_{ij}$ (11)
where $h_{i}=h(a_{i},w_{i},x_{i})$ and $k_{ij}=k((a_{i},z_{i},x_{i}),(a_{j},z_{j},x_{j}))$.
To prevent overfitting, we add an L2 penalty $\Lambda[h,\theta_{h}]=\sum_{i}\theta_{h,i}^{2}$, where $\theta_{h}$ are the model parameters. The penalized risk functions are:
$\hat{R}_{k,U,\lambda,n}(h)=\hat{R}_{k,U,n}(h)+\lambda\Lambda[h,\theta_{h}]$ (12)
$\hat{R}_{k,V,\lambda,n}(h)=\hat{R}_{k,V,n}(h)+\lambda\Lambda[h,\theta_{h}]$ (13)
The final loss function for training the neural networks is:
$\mathcal{L}_{proximal}=(Y-h(A,W,X))^{T}K(Y-h(A,W,X))+\lambda\Lambda[h,\theta_{h}]$ (14)
where $(Y-h(A,W,X))$ is a vector of residuals from the neural network’s predictions and $K$ is a kernel matrix with entries $k_{ij}$. We use an RBF kernel for $k$. For the U-statistic variant (NMMR-U), we set the main diagonal of $K$ to zero, while for the V-statistic variant (NMMR-V), we include the main diagonal elements.